OpenStudy (anonymous):

Can someone help me factor this quadratic? 2x^2 - 7x + 5

OpenStudy (anonymous):

You can write it as, 2x^2 - 2x - 5x + 5 = 2x(x-1) -5(x-1) = (x-1) (2x-5) So the roots of this quadratic equation are x=1 and x= 5/2

OpenStudy (anonymous):

Nice !! Thanks. Could I have used the Quadratic Formula on this one?

OpenStudy (anonymous):

Yes but only if this is an equation i.e, 2x^2 -7x + 5 =0 Then only we can talk about using the quadratic formula and its factors... Hope it makes sense....

OpenStudy (anonymous):

Perfect. Thanks for your help.

OpenStudy (anonymous):

My pleasure.